# Non-Abelian anyons in a periodically-driven Abelian model

### 06 Feb 2024 v1

@@ Abstract
We show that non-Abelian anyons can emerge from an Abelian topologically-ordered model subject to time-periodic driving, with the specific example of Ising anyons in a driven toric-code model. The toric code possesses fermionic and bosonic quasiparticles which are mutual semions, namely they see each other as $\pi$ fluxes. The Floquet modulation addresses the fermionic quasiparticles only and is designed to induce a topologically non-trivial band structure, realizing drive-assisted $p+ip$ pairing in the high-frequency regime. As a result, the fermions fractionalize into Floquet-Majorana zero modes hosted by the bosonic quasiparticles, which then develop non-Abelian character. These properties are analyzed through high-frequency expansions in a quasiparticle picture and, numerically, via computation of the exact quasienergy spectra and of the non-Abelian exchange phases. Our findings shed light on the nonequilibrium physics of driven topologically-ordered quantum matter and may facilitate the observation of non-Abelian behaviour in engineered quantum systems.
@@
Quasiparticles with anyon quantum statistics are predicted to appear as gapped excitations in topologically-ordered quantum phases of matter, such as fractional quantum Hall and spin liquids [1, 2]. Braiding anyons of Abelian type can change the many-body wavefunction by a phase factor. Braiding non-Abelian anyons, instead, can induce a unitary transformation in a topologically-protected degenerate ground-state subspace, enabling fault-tolerant quantum computation [3, 1]. Although realizing and controlling non-Abelian topological order in experiments remains elusive, the observation of Abelian anyons has been recently attained in engineered quantum systems, ranging from ultracold atoms in optical lattices [4, 5] to Rydberg atom arrays [6] and superconducting circuits [7]. This progress strongly motivates the study of topologically-ordered phases out of equilibrium, such as subject to external control, which may both provide a stepping stone towards non-Abelian physics [8, 9, 10, 11] and unveil novel effects, with first examples given by radical chiral Floquet phases [12, 13] and fractionalized prethermalization [14].
@@
Here we show that time-periodic driving can effectively turn an Abelian topologically-ordered model into a non-Abelian one. This is demonstrated with the example of Kitaev’s toric code [3], the paradigmatic $\mathds{Z}_{2}$ spin liquid [2]: under Floquet high-frequency modulation, the system develops non-Abelian Ising anyons, otherwise predicted in $p+ip$ topological superconductors [15], Kitaev’s honeycomb model (with a Haldane mass) [8] and the Pfaffian fractional-quantum-Hall state [16]. The toric code hosts bosonic ($\mathbb{Z}_{2}$ vortex) and fermionic ($\mathbb{Z}_{2}$ spinon) quasiparticle excitations behaving like mutual semions, namely seeing each other as $\pi$ fluxes. We construct a local Floquet drive, addressing the fermion quasiparticles, which attributes them a topological band structure in the high-frequency regime, effectively forcing them into a $p+ip$ superconducting phase. In the topological phase, attained at appropriate driving amplitude and frequency, the fermions fractionalize into Floquet-Majorana modes localized at the bosonic $\mathbb{Z}_{2}$ vortices, which then acquire non-Abelian character. This is substantiated by analytical analyses based on high-frequency expansions and numerical determination of the quasienergy spectra and of the vortex exchange phases. Our results shed light on the properties of nonequilibrium interacting topological phases and establish Floquet driving as a potential bridge between them. Moreover, they pave the way towards the development of novel protocols to Floquet-engineer non-Abelian anyons building on the control of Abelian phases recently attained [6] or predicted [17, 18, 19].
@@
The driven toric code model. The Abelian toric-code model [3, 20] describes two-level systems (spins) situated on the bonds of a square $L\times L$ lattice [Fig. 1(a)], with $L$ even hereafter, and ruled by the Hamiltonian
@@
The labels $p$ and $v$ denote plaquettes and vertices on the lattice, $g$ is a coupling constant, $\hat{A}_{v}=\prod_{i\in v}\hat{X}_{i}$ and $\hat{B}_{p}=\prod_{i\in p}\hat{Z}_{i}$ are four-spin interactions and $\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}$ are Pauli matrices pertaining to the $i$th spin. When convenient, we will also interpret $p=(p_{x},p_{y})$ and $v=(v_{x},v_{y})$ as coordinates of two separate plaquette and vertex lattices, respectively, both with lattice vectors $x=(1,0)$ and $y=(0,1)$. All four-spin operators composing $\hat{H}_{\rm tc}$ commute with each other, $[\hat{A}_{v},\hat{A}_{v^{\prime}}]=[\hat{A}_{v},\hat{B}_{p}]=[\hat{B}_{p},\hat{B}_{p^{\prime}}]=0$, making the model exactly solvable. Since $\hat{B}^{2}_{p}=\hat{A}^{2}_{v}=\mathds{1}$, the eigenstates can be labelled by the eigenvalues $A_{v}=\pm 1$ and $B_{p}=\pm 1$ for each vertex and plaquette. On a torus, where it can be shown that $\prod_{v}\hat{A}_{v}=\prod_{p}\hat{B}_{p}=\mathds{1}$, the eigenstates are further divided into four superselection sectors, identified by the ±$\pm 1$ eigenvalues of two operators $\hat{O}_{x}=\prod_{i\in\Lambda_{x}}{\hat{X}_{i}}$ and $\hat{O}_{y}=\prod_{i\in\Lambda_{y}}{\hat{X}_{i}}$ spanning the torus along horizontal ($\Lambda_{x}$) and vertical ($\Lambda_{y}$) noncontractible circles, respectively, and commuting with all $\hat{A}_{v}$ and $\hat{B}_{p}$ [Fig. 1(a)]. The model has four ground states, distinguished by the eigenvalues of $\hat{O}_{x}$ and $\hat{O}_{y}$, which satisfy $A_{v}=B_{p}=1$. Flipping one value $A_{v}=-1$ ($B_{p}=-1$), costs an excitation energy $g$ and is interpreted as the creation of a quasiparticle, called $e$ ($m$), on top of the ground state ($\mathds{1}_{\rm tc}$), which is localized at the corresponding vertex $v$ (plaquette $p$). While $e$ and $m$ are (hardcore) bosons under self-exchange, they are mutual semions: the wavefunction picks up a $\pi$ phase when $m$ is wound around $e$, or vice versa. This implies that a bound state $\psi=e\times m$ of an $e$ and $m$ particle, costing an energy $\mu_{\psi}=2g$, is a fermion, while still being a semion with respect to a separate $e$ or $m$. These quasiparticles, including $\mathds{1}_{\rm tc}$, represent the four anyon types in the toric code. They are Abelian, since their non-trivial fusion has only one possible outcome: $\quad e\times m=\psi\ ,\quad e\times\psi=m\ ,\quad m\times\psi=e\ .$
@@
We add to $\hat{H}_{\rm tc}$ a local time-periodic drive, whose goal is to induce motion of $\psi$ fermions and imbue them with a topologically non-trivial band structure. As detailed in the following, non-Abelian character will emerge thanks to the key interplay between the fermion band topology, achieved through Floquet engineering, and the semionic relation between $\psi$ and $e$ Abelian anyons, provided by the underlying toric-code model. The drive Hamiltonian is chosen as
$\hat{H}_{d}(t)=-\sum_{p}\big{[}d_{x}(t)\hat{X}_{p,R}\hat{Z}_{p,B}+d_{y}(t)\hat{X}_{p,T}\hat{Z}_{p,L}\big{]},$ (2)
@@
where the subscripts $({p,j})$, with $j\in\{B,R,T,L\}$ denote the spins at the bottom, right, top or left of plaquette $p$, respectively [Fig. 1(a)]. Two-body couplings of the form $\hat{X}_{i}\hat{Z}_{j}$ on adjacent spins induce $\psi$ movement and pairing, as sketched in Fig. 1(a): they consistently displace a joint pair of neighbouring $e$ and $m$, as well as create and annihilate pairs of them. Indeed, since $\hat{X}_{i}$ anticommutes with $\hat{B}_{p}$ on neighbouring plaquettes $p$ and $p^{\prime}$ and commutes with $\hat{A}_{v}$, its action on an eigenstate flips the corresponding eigenvalues $B_{p}$ and $B_{p^{\prime}}$, while preserving the $e$-anyon parity $(-1)^{B_{p}+B_{p^{\prime}}}$. Hence, if a $m$-particle composing a fermion is present at $p$, it will tunnel to $p^{\prime}$ (and vice versa). If no fermion or a pair is present, the associated pair of $m$ will be created or annihilated, respectively. An equivalent analysis applies to $e$-motion on nearby vertices induced by $\hat{Z}_{j}$. The time-periodic control functions $d_{x}(t)=d_{x}(t+T)$ and $d_{y}(t)=d_{y}(t+T)$, to be specified below, govern processes along the horizontal and vertical direction of the lattice, respectively.
@@
Figure 1: (a) Spin lattice of the toric code. The $m$ anyons live on plaquettes $p$ (orange), while $e$ anyons live on vertices $v$ (blue), and their bound state forms a fermion $\psi$. The eigenvalues of the operators $\hat{O}_{x}$ and $\hat{O}_{y}$ define four superselection sectors in the undriven system. The operators $\hat{X}_{i}\hat{Z}_{j}$ depicted induce fermion tunnelling and pairing. Dashed lines indicate periodic boundaries. (b) Lattice for $\psi$ fermions resulting from the quasiparticle mapping. The drives $d_{x}$ and $d_{y}$ rule fermion tunnelling and pairing along the horizontal and vertical direction, respectively, with the latter including also a string operators $\hat{\mathcal{S}}_{p}^{(y)}$ reproducing the semionic statistics of $\psi$ and $e$. (c) Drive-assisted fermion pairing, restoring otherwise off-resonant ($J\ll\mu_{\psi}$) processes with rate $\Delta$ and phase $\phi$.
@@
To focus on the behaviour of the $\psi$ quasiparticles, we adopt the quasiparticle picture of Ref. [21] [details in the Supplemental Material (SM) [22]]. The spins’ Hilbert space is mapped to the tensor product of Fock spaces for $e$ hardcore-boson and $\psi$ fermion occupations, and of a four-dimensional Hilbert space reflecting four different superselection sectors. In a given sector, the spin operators map to creation and annihilation operators of bosons on vertices ($\hat{b}_{v}^{\dagger},\hat{b}_{v}$) and of fermions on plaquettes ($\hat{f}_{p}^{\dagger},\hat{f}_{p}$), with the toric-code ground state |$\ket{0_{\rm tc}}$ representing the quasiparticle vacuum. The Hamiltonian $\hat{H}_{\rm tc}$, up to constant energy shifts, maps to $\hat{H}_{e}+\hat{H}_{e\psi}^{(0)}$ with the Hamiltonians $\hat{H}_{e}=g\sum_{v}\hat{b}_{v}^{\dagger}\hat{b}_{v}$ and $\hat{H}_{e\psi}^{(0)}=g\sum_{p}[1-\hat{b}^{\dagger}_{v(p)}\hat{b}_{v(p)}]\hat{f}_{p}^{\dagger}\hat{f}_{p}$, describing $e$ and $\psi$ particles at rest, respectively, where $v(p)$ denotes the vertex to the bottom-left of $p$. The drive $\hat{H}_{d}(t)$ maps to a coupling Hamiltonian $\hat{H}_{e\psi}^{(d)}(t)$, which describes the driven fermion tunnelling and pairing as well as anyon effects resulting from the fermion motion with respect to the $e$ bosons. Since $\hat{H}_{d}(t)$ drives only the fermions, one can show that $[\hat{H}_{e},\hat{H}_{e\psi}^{(d)}(t)]=0$ and analyse separately the fermion Hamiltonian $\hat{H}_{\psi}(t)=\bra{\vec{n}^{e}}\hat{H}_{e\psi}^{(0)}+\hat{H}_{e\psi}^{(d)}(t)\ket{\vec{n}^{e}}$ corresponding to a fixed ground state distribution of $e$ particles on the lattice, encoded in a $e$-boson Fock state |$\ket{\vec{n}^{e}}=(\hat{b}_{v_{1}}^{\dagger})^{n_{1}}\dots(\hat{b}_{v_{N}}^{\dagger})^{n_{N}}\ket{0_{\rm tc}}$. The presence of $e$-bosons at vertices $\{v\}$ is assumed to derive from preparing the ground state of a toric-code Hamiltonian (1) with inverted coupling $g\to\tilde{g}=-g$ for the vertices $\{v\}$. Then, the fermion Hamiltonian becomes
@@
While ${\mathcal{S}}_{p}^{(x)}=1$ in Eq. (3), the string ${\mathcal{S}}_{p}^{(y)}=\bra{\vec{n}^{e}}\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}\ket{\vec{n}^{e}}$ explicitly enforces the semionic statistics between $\psi$ and $e$. Here, $\mathrm{R}[v(p)]$ denotes all the vertices to the right of the spin shared by the plaquettes $p$ and $p+y$ [as in Fig. 1(b)]. Considering a fermion tunnelling on a loop encircling a $e$ particle, ${\mathcal{S}}_{p}^{(r)}$ indeed reproduces the accumulation of $\pi$ phase.
Floquet engineering. The driven Hamiltonian (3) describes spinless superconducting fermions, with tunnelling and pairing induced by $\hat{H}_{d}(t)$. They are further coupled to a deconfined $\mathbb{Z}_{2}$ gauge field given by a background distribution of $\mathbb{Z}_{2}$ vortices ($e$ particles), through $\hat{\mathcal{S}}_{p}^{(r)}$. These ingredients suggest a potential analogy to $p+ip$ topological superconductors [15, 23, 1], if the bulk fermion bands are gapped and topological. In such systems, fermions can fractionalize into Majorana zero modes localized at the $\mathbb{Z}_{2}$ vortices, which then become non-Abelian Ising anyons with non-trivial fusion rules $e\times e=\mathds{1}_{\rm is}+\psi$, $\psi\times\psi=\mathds{1}_{\rm is}$, $\psi\times e=e$ [1, 8]. The multiple fusion channels for $e$, yielding either the vacuum $\mathds{1}_{\rm is}$ or a fermion $\psi$, qualify them as non-Abelian. In the lattice model (3), $p+ip$ symmetry requires inequivalent pairing terms $\Delta_{r}f_{p}^{\dagger}f_{p+r}^{\dagger}$ along $x$ and $y$ of the form $(\Delta_{x},\Delta_{y})=(\Delta,i\Delta)$ [24]. However, since only control of the underlying toric-code spins is permitted, rather than on the quasiparticle pairing terms directly, no choice of constant $d_{x}$ and $d_{y}$ can realize the desired $p+ip$ coupling. Indeed, the functions $d_{r}(t)$ need be real-valued and they induce tunnelling and pairing with equal real rate. To overcome these limitations, we employ a periodic time-dependence in $d_{r}(t)$ and Floquet engineering [25, 26]. The driving functions are chosen of the form
$d_{r}(t)=J+2\Delta\cos(\omega t+\phi_{r}).$ (4)
@@
$\hat{H}_{F}=$ $-\mu\sum_{p}\hat{f}_{p}^{\dagger}\hat{f}_{p}$
-- $\sum_{r=x,y}{\mathcal{S}}_{p}^{(r)}\ \big{(}\Delta e^{i\phi_{r}}\hat{f}_{p}\hat{f}_{p+r}+J\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.})\ .$ (5)
@@
The stroboscopic propagator reads $\hat{U}(T)=\hat{R}(T)e^{-i\hat{H}_{F}T}$. At time $T$, the transformation $\hat{R}(t)$ reduces to the total fermion parity of the system, $\hat{R}(T)=\exp(-i\pi\sum_{p}\hat{f}^{\dagger}_{p}\hat{f}_{p})=(-1)^{\sum_{p}\hat{f}^{\dagger}_{p}\hat{f}_{p}}$, which is a conserved quantity in the driven (and thus also the Floquet) Hamiltonian. Since quasiparticles in the toric code can only be created in pairs, only even-parity states are physical and, given that $R(T)=1$ in their subspace, their stroboscopic dynamics is ruled by $\hat{H}_{F}$ of Eq. (5).
From Eq. (5), we obtain pairing with $p+ip$ symmetry by choosing a circular-like shaking, with phase delay $\phi_{y}-\phi_{x}=\pi/2$ between horizontal and vertical drives. The high-frequency effective Hamiltonian, Eq. (5), then describes spinless superconducting $\psi$ fermions with $p+ip$ pairing, a chemical potential $\mu$ controlled by the detuning of the drive from the fermion-pair chemical potential of the toric code and coupled to the $\mathbb{Z}_{2}$ vortices via $\hat{\mathcal{S}}_{p}^{(r)}$. We next verify that the driven model hosts Floquet-Majorana modes and realizes Ising anyons.
@@
Figure 2: (a) Splitting between fermion ground states for $\mu=-2J$ in different superselection sectors, characterized by periodic ($P_{x/y}$) or anti-periodic ($A_{x/y}$) boundary conditions and having energy $E_{\rm GS}^{(x,y)}$ [$(x,y)$ denotes the boundary condition]. The $P_{x}P_{y}$ ground state has odd parity. Solid lines and circles denote results in the high-frequency regime ($J=\Delta=0.1\mu_{\psi}$), while dashed lines and squares correspond to the anomalous phase ($J=\Delta=0.15\mu_{\psi}$) (b) FBdG quasienergy spectrum and edge modes on a $L_{x}\times L_{y}=20\times 200$ cylinder, periodic along $x$, for $J=\Delta=0.1\mu_{\psi}$. The color scale indicates the inverse participation ratio of the corresponding FBdG eigenstate. (c) FBdG quasienergy spectrum for $\mu=-2J$ and $L=24$ for increasing $J=\Delta$. (d) Coefficients $\log_{10}|u_{p}^{T}|^{2}$ of the zero mode defined by $\hat{\alpha}_{1}^{\dagger}=\sum_{p}(v_{p}\hat{f}_{p}+u_{p}^{T}\hat{f}_{p}^{\dagger})$ with $v_{p}\simeq u_{p}^{*}$, for $J=\Delta=0.1\mu_{\psi}$ and $L=24$.

### 28 Feb 2024 v2

@@ Abstract
We show that non-Abelian anyons can emerge from an Abelian topologically-ordered model subject to local time-periodic driving, with the specific example of Ising anyons in a driven toric-code model. The quasiparticle excitations of the toric code, realizing an Abelian-anyon theory, include fermionic and bosonic quasiparticles which see each other as $\pi$ fluxes, namely they result in the accumulation of a $\pi$ phase if wound around each other. The emergence of non-Abelian behaviour is explained in the fact that the modulation can engineer a band topology for the fermionic quasiparticles, by Floquet engineering complex phases of fermion pairing in the high-frequency regime. This induces the fractionalization of the fermions into Floquet-Majorana modes bound to the bosons. The latter then develop non-Abelian anyon character akin to vortices in a topological superconductor. Our findings shed light on the nonequilibrium physics of driven topologically-ordered quantum matter and may facilitate the observation of non-Abelian behaviour in engineered quantum systems.
@@
Quasiparticles with anyon quantum statistics are predicted to appear as gapped excitations in topologically-ordered quantum phases of matter, such as fractional quantum Hall and spin liquids [1, 2]. Braiding anyons of Abelian type can change the many-body wavefunction by a phase factor. Braiding non-Abelian anyons, instead, can induce a unitary transformation in a topologically-protected degenerate ground state, enabling fault-tolerant quantum computation [3, 1]. Although realizing and controlling non-Abelian topological order remains elusive, Abelian anyons have been recently observed in engineered quantum systems, ranging from ultracold atoms in optical lattices [4, 5] to Rydberg-atom arrays [6] and superconducting circuits [7]. This progress motivates the study of topologically-ordered phases out of equilibrium, such as subject to external control, which may both provide a stepping stone towards non-Abelian physics [8, 9, 10, 11, 12] and unveil novel effects, with examples given by radical chiral Floquet phases [13, 14] and fractionalized prethermalization [15].
@@
We show that time-periodic driving can effectively turn an Abelian topologically-ordered model into a non-Abelian one. This is demonstrated with the example of Kitaev’s toric code [3], the paradigmatic $\mathds{Z}_{2}$ quantum spin liquid [2]. Abelian anyons in the toric code emerge as bosonic and fermionic quasiparticle excitations which see each other as $\pi$ fluxes: Winding a fermion around a boson (and vice versa) results, unusually, in the accumulation of a $\pi$ Peierls-like phase. The Floquet modulation of the physical spins fundamentally alters this picture by driving one Abelian-anyon type, the fermion, into a phase characterized by a non-trivial band topology. This is achieved by Floquet-engineering fermion pairing with a complex phase, mimicking key features of topological superconductors [16, 1]. Such processes induce the fractionalization of the fermionic quasiparticles into pairs of Majorana modes binding to the other (undriven) anyon type, namely the bosonic quasiparticles. We verify that these Majorana-carrying bosons behave like non-Abelian Ising anyons, known in other contexts such as $p+ip$ superconductors [16], Kitaev’s honeycomb model [8] and the Pfaffian quantum Hall state [17]. Our analysis involves analytical high-frequency expansions and the numerical determination of quasienergy spectra and of the non-Abelian exchange phases. These results shed light on the properties of nonequilibrium interacting topological phases and establish Floquet driving as a potential link between them. Moreover, they pave the way towards the development of novel protocols to Floquet-engineer non-Abelian anyons [18, 19] building on the control of Abelian phases recently attained [6] or predicted [20, 21, 22].
@@
The driven model. The Abelian toric-code model [3, 23] describes two-level systems (spins) on the bonds of a square $L\times L$ lattice [Fig. 1(a)], with $L$ even hereafter, and ruled by the Hamiltonian
@@
The labels $p$ and $v$ denote lattice plaquettes and vertices, $g$ is a coupling constant, $\hat{A}_{v}=\prod_{i\in v}\hat{X}_{i}$ and $\hat{B}_{p}=\prod_{i\in p}\hat{Z}_{i}$ are four-spin interactions and $\hat{X}_{i},\hat{Y}_{i},\hat{Z}_{i}$ are Pauli matrices for the $i$th spin. When convenient, we will also interpret $p=(p_{x},p_{y})$ and $v=(v_{x},v_{y})$ as coordinates of two separate plaquette and vertex lattices, respectively, both with lattice vectors $x=(1,0)$ and $y=(0,1)$. All four-spin operators composing $\hat{H}_{\rm tc}$ commute with each other, $[\hat{A}_{v},\hat{A}_{v^{\prime}}]=[\hat{A}_{v},\hat{B}_{p}]=[\hat{B}_{p},\hat{B}_{p^{\prime}}]=0$, making the model exactly solvable. Since $\hat{B}^{2}_{p}=\hat{A}^{2}_{v}=\mathds{1}$, the eigenstates can be labelled by the eigenvalues $A_{v}=\pm 1$ and $B_{p}=\pm 1$ for each vertex and plaquette. On a torus, where it can be shown that $\prod_{v}\hat{A}_{v}=\prod_{p}\hat{B}_{p}=\mathds{1}$, the eigenstates are further divided into four superselection sectors, identified by the ±$\pm 1$ eigenvalues of two operators $\hat{O}_{x}=\prod_{i\in\Lambda_{x}}{\hat{X}_{i}}$ and $\hat{O}_{y}=\prod_{i\in\Lambda_{y}}{\hat{X}_{i}}$ spanning the torus along horizontal ($\Lambda_{x}$) and vertical ($\Lambda_{y}$) noncontractible loops and commuting with all $\hat{A}_{v}$ and $\hat{B}_{p}$ [Fig. 1(a)]. The model has four degenerate ground states, satisfying $A_{v}=B_{p}=1$, which are distinguished by the eigenvalues of $\hat{O}_{x}$ and $\hat{O}_{y}$. Flipping one eigenvalue, $A_{v}=-1$ ($B_{p}=-1$), costs an excitation energy $g$ and is interpreted as the creation of a pointlike quasiparticle $e$ ($m$) on top of the ground state $\mathds{1}_{\rm tc}$, which occupies the corresponding vertex $v$ (plaquette $p$). While $e$ and $m$ are hardcore bosons under self-exchange, they are mutual semions: the wavefunction picks up a Peierls-like $\pi$ phase when $m$ is wound around $e$, or vice versa. A bound state $\psi=e\times m$ of an $e$ and $m$ particle, costing an energy $\mu_{\psi}=2g$, is then a fermion, while still being a semion with respect to a separate $e$ or $m$. These quasiparticles, including $\mathds{1}_{\rm tc}$, represent the four anyon types in the toric code. They are Abelian, since their non-trivial fusion has a single possible outcome: $e\times m=\psi$, $e\times\psi=m$, $m\times\psi=e.$
@@
We add to $\hat{H}_{\rm tc}$ a drive in order to set the $\psi$-fermions in motion and imbue them with a topological band structure. Non-Abelian character will emerge from the key interplay between the fermion band topology, achieved through Floquet engineering, and the semion relation between the $\psi$ and $e$ Abelian anyons, provided by the underlying toric code. The drive is chosen as
$\hat{H}_{d}(t)=-\sum_{p}\big{[}d_{x}(t)\hat{X}_{p,R}\hat{Z}_{p,B}+d_{y}(t)\hat{X}_{p,T}\hat{Z}_{p,L}\big{]}.$ (2)
@@
The subscripts $({p,j})$, with $j\in\{B,R,T,L\}$ denote the spins at the bottom, right, top or left of $p$, respectively [Fig. 1(a)]. Two-body couplings of the form $\hat{X}_{i}\hat{Z}_{j}$ on adjacent spins induce $\psi$ tunnelling and pairing, as sketched in Fig. 1(a): they displace a joint pair of neighbouring $e$ and $m$ composing $\psi$, as well as create and annihilate pairs of them. Indeed, since $\hat{X}_{i}$ anticommutes with $\hat{B}_{p}$ on neighbouring plaquettes $p$ and $p^{\prime}$ while commuting with $\hat{A}_{v}$, its action on an eigenstate flips the eigenvalues $B_{p}$ and $B_{p^{\prime}}$, while preserving the joint $m$-anyon parity $(-1)^{B_{p}+B_{p^{\prime}}}$. Hence, if a $m$-particle composing a fermion is present at $p$, it tunnels to $p^{\prime}$ (and vice versa). If no fermion or a pair is present, the associated pair of $m$ will be created or annihilated. An equivalent analysis applies to $e$-motion on nearby vertices induced by $\hat{Z}_{j}$. The control functions $d_{x}(t)$ and $d_{y}(t)$ govern processes along the horizontal and vertical direction of the lattice, respectively.
@@
Figure 1: (a) Toric-code spin lattice. The $m$ anyons live on plaquettes $p$ (orange), while $e$ anyons live on vertices $v$ (blue), and their bound state forms a fermion $\psi$. The eigenvalues of $\hat{O}_{x}$ and $\hat{O}_{y}$ define four superselection sectors in the undriven system. The operators $\hat{X}_{i}\hat{Z}_{j}$ depicted induce fermion tunnelling and pairing. Dashed lines indicate periodic boundaries. (b) Lattice for $\psi$ fermions in the quasiparticle mapping. The controls $d_{x}$ and $d_{y}$ drive fermion tunnelling and pairing. Vertical processes include a boson operator $\hat{\mathcal{S}}_{p}^{(y)}$ reproducing the mutual statistics of $\psi$ and $e$. (c) Drive-assisted fermion pairing, restoring otherwise off-resonant ($J\ll\mu_{\psi}$) processes with rate $\Delta$ and phase $\phi_{r}$.
@@
Quasiparticle picture. To discuss the impact of the drive on the $\psi$ quasiparticles, we adopt the quasiparticle picture of Ref. [24] (details are given in the Supplemental Material (SM) [25]). The spins’ Hilbert space is mapped to the tensor product of Fock spaces for $e$ hardcore-boson and $\psi$ fermion occupations, and of a four-dimensional Hilbert space reflecting four superselection sectors. The spin operators map to creation and annihilation operators of bosons on vertices ($\hat{b}_{v}^{\dagger},\hat{b}_{v}$) and of fermions on plaquettes ($\hat{f}_{p}^{\dagger},\hat{f}_{p}$), with the toric-code ground state |$\ket{0_{\rm tc}}$ representing the quasiparticle vacuum. The Hamiltonian $\hat{H}_{\rm tc}$ maps to $\hat{H}_{e}+\hat{H}_{e\psi}^{(0)}$ with $\hat{H}_{e}=g\sum_{v}\hat{b}_{v}^{\dagger}\hat{b}_{v}$ and $\hat{H}_{e\psi}^{(0)}=2g\sum_{p}[1-\hat{b}^{\dagger}_{v(p)}\hat{b}_{v(p)}]\hat{f}_{p}^{\dagger}\hat{f}_{p}$, describing quasiparticles at rest [$v(p)$ denotes the vertex to the bottom-left of $p$]. The drive $\hat{H}_{d}(t)$ maps to $\hat{H}_{e\psi}^{(d)}(t)=-\sum_{p,r\in\{x,y\}}d_{r}(t)\hat{\mathcal{S}}_{p}^{(r)}(\hat{f}_{p}\hat{f}_{p+r}+\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.})$, describing fermion tunnelling and pairing but also containing a coupling to the $e$-bosons through $\hat{\mathcal{S}}_{p}^{(x)}=1$ and $\hat{\mathcal{S}}_{p}^{(y)}=\prod_{v\in{\mathrm{R}[v(p)]}}(-1)^{\hat{b}_{v}^{\dagger}\hat{b}_{v}}$. Here, $\mathrm{R}[v(p)]$ denotes the vertices to the right of the spin shared by the plaquettes $p$ and $p+y$, see Fig. 1(b). This coupling explicitly reproduces the accumulation of Peierls-like phase as the anyons move with respect to each other: Considering a fermion tunnelling on a loop encircling a $e$ particle, $\hat{\mathcal{S}}_{p}^{(r)}$ indeed yields the accumulation of $\pi$ phase. Since the drive does not induce boson motion, all bosonic terms are diagonal in the Fock basis |$\ket{\vec{n}^{e}}=(\hat{b}_{v_{1}}^{\dagger})^{n_{1}}\dots(\hat{b}_{v_{N}}^{\dagger})^{n_{N}}\ket{0_{\rm tc}}$ describing $e$ occupations. We can then analyze the fermion Hamiltonian $\hat{H}_{\psi}(t)=\bra{\vec{n}^{e}}\hat{H}_{e\psi}^{(0)}+\hat{H}_{e\psi}^{(d)}(t)\ket{\vec{n}^{e}}$ corresponding to a fixed distribution of $e$ particles. The presence of $e$-bosons at vertices $\{v\}$ is assumed to derive from preparing the ground state of a slightly modified Hamiltonian (1) with inverted coupling $g\to\tilde{g}=-g$ for the vertices $\{v\}$. The Hamiltonian $\hat{H}_{\psi}(t)$ then reads
@@
It describes spinless superconducting fermions coupled, through $\hat{\mathcal{S}}_{p}^{(r)}$, to an effective flux given by a background distribution of $e$ particles (also interpretable as $\mathbb{Z}_{2}$ vortices of a deconfined $\mathbb{Z}_{2}$ gauge field [2, 26]).
Floquet engineering. The above ingredients suggest a potential analogy to topological superconductors [16, 1, 27]. In such systems, non-Abelian anyons are predicted to emerge as vortices ($\sigma$) carrying a Majorana zero-mode [28]. Majorana modes attached to topological defects result from the fractionalization of the superconducting fermions when their bulk bands are gapped and topological, as is the case for pairing in so-called $p+ip$ symmetry [1]. In a lattice model as in (3), this corresponds to a complex-valued pairing coupling $\Delta_{r}f_{p}^{\dagger}f_{p+r}^{\dagger}$ along $x$ and $y$ of the form $(\Delta_{x},\Delta_{y})=(\Delta,i\Delta)$ [26]. The resulting so-called Ising anyons $\sigma$, $\mathds{1}_{\rm is}$ and $\psi$ feature non-trivial fusion rules $\sigma\times\sigma=\mathds{1}_{\rm is}+\psi$, $\psi\times\psi=\mathds{1}_{\rm is}$, $\psi\times\sigma=\sigma$ [1, 8]. The multiple fusion channels for the vortex $\sigma$, yielding either the vacuum $\mathds{1}_{\rm is}$ or a fermion $\psi$, qualify them as non-Abelian. We will show that the driven toric code reproduces faithfully this physics and non-Abelian anyons, with $e$ particles behaving like $\sigma$ vortices, for appropriate driving functions that Floquet-engineer $p+ip$ fermion pairing.
Since the control can only address the physical spins, rather than the quasiparticle pairing terms directly, no choice of time-independent $d_{x}$ and $d_{y}$ in Eq. (2) can achieve the goal. Indeed, the functions $d_{r}(t)$ need be real-valued and they induce tunnelling and pairing with equal real rate. We overcome these limitations employing a periodic modulation $d_{r}(t)=d_{r}(t+T)$ and Floquet engineering [29, 30]. The driving functions are chosen as
$d_{r}(t)=J+2\Delta\cos(\omega t+\phi_{r}),$ (4)
@@
$\hat{H}_{F}=$ $-\sum_{p}\Big{[}\mu\hat{f}_{p}^{\dagger}\hat{f}_{p}$
$+$ $\sum_{r=x,y}{\mathcal{S}}_{p}^{(r)}\ \big{(}\Delta e^{i\phi_{r}}\hat{f}_{p}\hat{f}_{p+r}+J\hat{f}_{p}^{\dagger}\hat{f}_{p+r}+\mathrm{H.c.})\Big{]}\ .$ (5)
@@
At time $T$, the transformation $\hat{R}(t)$ reduces to the total fermion parity, $\hat{R}(T)=(-1)^{\sum_{p}\hat{f}^{\dagger}_{p}\hat{f}_{p}}$, which is conserved by the driven (and thus also the Floquet) Hamiltonian. Since quasiparticles in the toric code can only be created in pairs, only even-parity states are physical. Given that $R(T)=1$ in their subspace, their stroboscopic dynamics is ruled by $\hat{H}_{F}$ of Eq. (5). We obtain the desired $p+ip$ pairing in $\hat{H}_{F}$ by choosing a circular-like shaking, with phase delay $\phi_{y}-\phi_{x}=\pi/2$ between horizontal and vertical modulations. The detuning of the drive from the excitation energy of a fermion pair in the toric code controls the effective chemical potential $\mu$, and the fermions are coupled to the $e$ particles via ${\mathcal{S}}_{p}^{(r)}$.
@@
Figure 2: (a) Splitting between fermion ground states for $\mu=-2J$ in different superselection sectors, having energy $E_{\rm GS}^{(x,y)}$ [$(x,y)$ denotes the boundary conditions, either periodic ($P$) or antiperiodic ($A$) along $x$ and $y$] in the high-frequency regime ($J=\Delta=0.1\mu_{\psi}$, solid lines and circles) and in the anomalous phase ($J=\Delta=0.15\mu_{\psi}$, dashed lines and squares). (b) Floquet-BdG quasienergies on a $L_{x}\times L_{y}=20\times 200$ cylinder, for $J=\Delta=0.1\mu_{\psi}$, and (c) on a $(A,A)$-torus in the presence of a $e$ pair, for increasing driving amplitude ($\mu=-2J$, $L=24$). The color-scale indicates the inverse participation ratio of the corresponding eigenstate. (d) Paths $P$ and $P^{\prime}$ involved in the vortex-exchange protocol. (e) Vortex exchange phase $\theta_{P}-\theta_{P^{\prime}}$ in the $\mathds{1}_{\rm is}$ (bottom) and $\psi$ (top) sectors at $\omega=2(\mu_{\psi}-2J)$, for different driving amplitude and system size.